OpenStudy (anonymous):

please Help with Trig... solve for all degree solutions and x if 0 deg < x< 360 deg: (sq root 3 tan x)-(2 sin x tan x)=0

OpenStudy (anonymous):

\[\sqrt{3} \tan x - 2\sin x \tan x = 0 \iff \tan x\ \left( \sqrt{3} - 2\sin x\right) = 0\] \[\implies \tan x = 0 \text{ or } \sin x = \frac{\sqrt{3}}{2} \]

OpenStudy (anonymous):

oh, wow I understand

OpenStudy (anonymous):

Thank you very much!